Hello, I got this question wrong on a test, so I was just wondering what I did wrong. Thanks for the help!
Find the Maximum or minimum value of y=1/3x^2+x+1 and determine for which value of x it occurs
Thanks again

Answer by ankor@dixie-net.com(22740)

Find the Maximum or minimum value of y=1/3x^2+x+1 and determine for which value of x it occurs
:
you want to find the axis of symmetry. Using the form y = ax^2 + bx + c,
the axis of symmetry is found by: x = -b/(2a)
:
In your equation; a = 1/3; b = 1; c = 1
:
x = -1/(2*(1/3)
x = -1/(2/3)
x = -3/2 or -1.5
:
We know it is a minimum, because the coefficient of x^2 is positive
:
to find the vertex, substitute -1.5 in the original equation.